answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What does non probability mean?

It is likely to mean deterministic. It means that the outcome of an event is known and not subject to probability.

It is likely to mean deterministic. It means that the outcome of an event is known and not subject to probability.

It is likely to mean deterministic. It means that the outcome of an event is known and not subject to probability.

It is likely to mean deterministic. It means that the outcome of an event is known and not subject to probability.

What is the probability of getting difference 2 with two dice?

The total number of ways to obtain the certain difference is 36. The number of ways of obtaining the difference of 2 with two dice is 8. How? List the possibilities of obtaining the difference of 2 with two dice:
  1. 3 and 1
  2. 4 and 2
  3. 5 and 3
  4. 6 and 4


These are 4 pairs of numbers to obtain a difference 2. Make note that we can also have the difference of two with these numbers "reversed". Then, we have 8 ways of obtaining the difference of 2. Therefore, the probability of obtaining the difference of 2 with two dice is 8/36.

What is the probability of obtaining exactly four tails in five flips?

Assume the coin is fair, so there are equal amount of probabilities for the choices.

There are two possible choices for a flip of a fair coin - either a head or a tail. The probability of getting a head is ½. Similarly, the probability of getting a tail is ½.

Use Binomial to work out this problem. You should get:

(5 choose 4)(½)4(½).

  • (5 choose 4) indicates the total number of ways to obtain 4 tails in 5 flips.
  • (½)4 indicates the probability of obtaining 4 tails.
  • (½) indicates the probability of obtaining the remaining number of head.

Therefore, the probability is 5/32.

What the probability that Fred will get a black sweet if there 3 yellow sweets 5 green sweets 7 red sweets 4 purple sweets 1 black sweet?

Step 1: Count the total number of sweets in a "bag"

There are 3 yellow, 5 green, 7 red, 4 purple and 1 black sweets. Add all these values up to obtain 20 sweets total.

Step 2: Determine the probability of the given event

Since we want to determine the probability of selecting a black sweet out of a "bag", we have:

1 / 20 = number of black sweets / total number of sweets

How many ways can CARS be arranged?

4*3*2*1 = 24 ways.

4*3*2*1 = 24 ways.

4*3*2*1 = 24 ways.

4*3*2*1 = 24 ways.

What is the probability of getting a six when you roll a dice?

The answer depends on how often you roll it! For one roll it is 1/6 but the probability increases to a near certainty as you increase the number of rolls.

How many ways can you arrange 5 objects in straight line?

That would be 5x4x3x2x1 or 5! or 120 ways to arrange those objects in a line.

How many ways can you arrange 4 different cubes?

The cubes can be arranged in a line in 24 orders, and each cube can be arranged in 24 orientations so the answer is 24*24 = 576 ways.

Of course, the question does not require then to be in a line so there are really infinitely many ways: for example one cube in your house, one in your garden, one in my house and one in my garden. Move one of them a bit and that's a different arrangement.

What is the probability of getting a red heart in a deck?

The probability of drawing a red heart is 1 in 4. This is the same as the probability of drawing a heart, as red is included as a superset of hearts.

Standard deviation is denoted by?

The letter s, or by sd. The Greek lower case sigma is also used.

What is the probability of rolling either a 5 or 6 or 8 or 9 before rolling a 7 with two dice?

Consider the three events:

A = rolling 5, 6, 8 or 9.

B = rolling 7

C = rolling any other number.

Let P be the probability of these events in one roll of a pair of dice.

Then P(A) = P(5) + P(6) + P(8) + P(9) = 18/36 = 1/2

P(B) = P(7) = 6/36 = 1/6

and P(C) = 1 - [P(A) + P(B)] = 1/3

Now P(A before B) = P(A or C followed by A before B)

= P(A) + P(C)*P(A before B)

= 1/2 + 1/3*P(A before B)

That is, P(A before B) = 1/2 + 1/3*P(A before B)

or 2/3*P(A before B) = 1/2

so that P(A before B) = 1/2*3/2 = 3/4

What is the probability of a vowel being picked at random from the alphabet?

Probability is given as Desired Outcomes divided by total number of outcomes.

For the probability of picking a vowel, desired outcomes are : a,e,i,o,u (5)

Total no. of outcomes is the entire alphabet set from a to z (26)

Hence, the required probabilty is 5/26

What is the probability of getting a multiple of 2 when you roll a 6 sided die?

When you roll a 6 sided die there are 6 possible outcomes. 1,2,3,4,5,6. Out of these 3 are multiples of 2.

Therefore the Probability of getting a multiple of 2 is (3/6) which simplifies to (1/2) or a half.

What is the probability of choosing a queen or a king from a standard deck of cards?

The probability of drawing a queen or king, in a single randomly drawn card, is 2/13.

The probability of drawing one when you draw 45 cards without replacement is 1.

The probability of choosing has nothing t do with the probability of drawing the card. I can choose a king but fail to find one!

What is the probability of rolling a 4 on a number cube and flipping tails on a coin?

The probability of rolling any number on a cube can be represented by the formula: X / the number of variables. Since any cube has 6 sides, the probability of rolling any of the numbers 1 through 6 on the cube, can be represented by the formula: X = 1 / 6

= 16.66%

The odds or probability of flipping a coin and landing it on either side can be represented by X = the requested result / the number of variables

= 1 /2

= 50%

Therefore, given the two questions of probability, there is a much greater chance of landing a coin on "tails" rather than rolling a "4".

What is the mean and SD of a Z score?

Mean = 0

SD = 1

The whole point of converting to a Z-score is that you have a Standard Normal distribution ie a N(0, 1) distribution.

Trending Questions
A box contains 75 balls numbered from 1 to 75. If 12 balls are drawn with replacement what is the probability that at least two of them have the same number? What is the probability that your birthday will fall on Saturday and Sunday? Forty percent of all registered voters in a national election are female A random sample of 5 voters is selected The probability that the sample contains 2 female voters is? What is the probability of a head on the first flip and a tail on the second flip? What must you do to determine the probability of multiple independent events happening a certain way? Can a ace beat a joker in cards? What are the odds of rolling a 6 or an 8 before rolling a 7 in craps? What is the probability of picking two red crayons if you pick two in a row without replacing the first one from 11 crayons including 4 red 3 yellow and 4 green crayons? How many aces are there in 52 deck of card? The t distribution is used to construct confidence intervals for the population mean when the population standard deviation is unknown? What is the advantage of a one-tailed test? How many ways can 9 books be arranged on a shelf so that 5 of the books are always together? What is the probability of drawing two black kings from a standard deck of 52 cards? Which is better sampling designscientific or non-scientific? Range of probability? Two dice are thrown. What is the probability of getting the same number? What is the variance for the uniform random variable? Is the area under the standard normal distribution to the left of z equals 0 negative? What is the binomial probability if the mean value is 3 and the variance is 1.2? How many different ways can you arrange the number 7532?